# Spectral-Spatial Transformer with Active Transfer Learning for Hyperspectral Image Classification

### v1

@@ II-A3 Transformer Encoder Block
where the queries, keys, and values are computed as:
$\mathbf{Q}_{i}=\mathbf{Z}_{i}\mathbf{W}_{Q},\quad\mathbf{K}_{i}=\mathbf{Z}_{i}\mathbf{W}_{K},\quad\mathbf{V}_{i}=\mathbf{Z}_{i}\mathbf{W}_{V}$ (6)
The attention scores are calculated using the scaled dot-product attention:
@@
where $d_{k}$ is the dimension of the keys and $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, $\mathbf{W}_{V}$ are learnable weight matrices for the queries, keys, and values, respectively. The output of the multi-head attention layer is passed through a feedforward network consisting of two dense layers with a ReLU activation in between:
$\mathbf{F}(\mathbf{Z}_{i})=\max(0,\mathbf{Z}_{i}\mathbf{W}_{1}+b_{1})\mathbf{W}_{2}+b_{2}$ (8)
The final output of the transformer encoder layer is given by:
@@
where LayerNorm represents layer normalization and Dropout is applied for regularization.
II-A4 Cross-Attention Mechanism
The SST model integrates a cross-attention mechanism to enhance feature representation across patches. Let $\mathbf{Q}_{c}\in\mathbb{R}^{d}$ be a learnable query token and $\mathbf{K}_{c},\mathbf{V}_{c}$ be the keys and values derived from the patch embeddings. The cross-attention is computed as:
@@ II-A4 Cross-Attention Mechanism
This allows the model to incorporate global context from all patches while maintaining spatial-spectral coherence.
II-A5 Classification Head
The output of the transformer encoder and cross-attention mechanisms are concatenated, followed by two fully connected layers with ReLU activations:
@@ II-A5 Classification Head
where $\mathbf{O}_{\text{final}}\in\mathbb{R}^{C}$ is the final class prediction, and $C$ is the number of target classes.
II-B Active Learning
We employ an active learning technique to iteratively query the most informative samples from a pool of data. The following query strategies are applied:
II-B1 Hybrid Query Strategy
The hybrid query strategy combines uncertainty and diversity sampling. Let $\mathcal{U}$ and $\mathcal{D}$ denote the set of samples selected by uncertainty and diversity sampling, respectively. The final query set $\mathcal{Q}$ is formed as:
$\mathcal{Q}=\mathcal{U}(\mathbf{x}_{i})\cup\mathcal{D}(\mathbf{x}_{i})$ (14)
where $\mathcal{Q}$ ensures that both uncertain and diverse samples are included in the labeled set. The hybrid strategy combines uncertainty and diversity sampling to balance exploration and exploitation in the active learning setting. Uncertainty sampling selects samples for which the model is least confident. For a sample $\mathbf{x}_{i}$, the uncertainty score is calculated as:
@@ II-B1 Hybrid Query Strategy
where $p(y|\mathbf{x}_{i})$ is the predicted probability of the class for $\mathbf{x}_{i}$. Samples with the highest uncertainty scores are prioritized for labeling. Spatial-spectral neighborhood diversity sampling ensures that the selected pixels from the HSI are representative of diverse spatial-spectral characteristics within their local neighborhoods. This approach computes diversity based on the average pairwise spectral distance among neighboring pixels. Mathematically, this can be framed as:
@@
where the diversity of a set $S$ of pixels is defined as:
$\text{Diversity}(S)=\frac{1}{m(m-1)}\sum_{j=1}^{m}\sum_{k=1,k\neq j}^{m}d_{jk}$ (17)
where $d_{jk}$ is the pairwise Euclidean distance between the spectral features of pixels $j$ and $k$ within their respective neighborhoods, and $m$ is the number of pixels in the local neighborhood defined as:
$m=(n_{\text{neighborhood}})^{2}$ (18)
The overall diversity for each pixel is computed as:
@@
where $(h,w)$ represents the coordinates of the pixel in the HSI. The query indices of the most diverse pixels are selected based on:
@@
This ensures that the final selection of pixels captures the spatial-spectral diversity, enhancing the effectiveness of AL strategies.
II-C Model Training and Evaluation
The SST model is trained using a categorical cross-entropy loss function:
@@ II-C Model Training and Evaluation
where $y_{ic}$ is the ground truth label and $\hat{y}_{ic}$ is the predicted probability for class $c$ for the $i$-th sample. The model is optimized using Adam with an initial learning rate of $1\times 10^{-4}$. We evaluate the model on various metrics including overall accuracy (OA), average accuracy (AA), per-class accuracy (PA), and the kappa coefficient ($\kappa$). The kappa coefficient is calculated as:
@@
where $p_{o}$ is the observed agreement and $p_{e}$ is the expected agreement by chance.
III Experimental Settings
The experimental setup for the SST-ATL incorporates several key parameters and methodologies to ensure clarity and reproducibility. Following data (HSI cube and ground truths) reading, image cubes are created from the processed data, and the dataset is split into training, validation, and test sets with ratios defined as training ratio set to 0.01%, pool ratio 0.49%, and test ration 0.50%. This results in an initial training sample count of 1% of the dataset, while the pool set comprises 49% of the remaining samples and the test set is 50% of the original data.
@@ III Experimental Settings
The model training is executed with several defined hyperparameters. The learning rate is set at 0.001 with a decay of $1\times 10^{-6}$, utilizing the Adam optimizer. The training process spans 50 epochs with a batch size of 56. The architecture of SST incorporates multiple transformer layers and attention mechanisms. Specifically, the model includes 4 transformer layers, with each layer employing 8 attention heads. The dimensionality of the model ($d_{\text{model}})$ is set to 54, and the feedforward network’s dimensionality is $4\times 64=256$. The dropout rate is fixed at 0.1 to prevent overfitting, and the layer normalization epsilon value is set to $1\times 10^{-6}$.
The model’s input shape is defined as ($W\times W\times k$), where $W$ represents the patch size, which is derived from the dimensions ($k$) of the input data. The model employs a patch embedding mechanism to transform input data into manageable segments for processing within the transformer architecture. The embeddings are enhanced with positional encoding to preserve spatial information, which is essential in processing sequential data in transformer networks.
The experimental framework incorporates AL techniques. Initially, the model is trained on a small dataset, after which AL iterations take place. The hybrid query strategy is utilized to query new samples from the pool, combining uncertainty sampling and diversity sampling methods. The queried samples are integrated into the training set, and the model undergoes fine-tuning through retraining with a subset of the last layers frozen. This iterative process continues for several cycles, ensuring the model adapts and improves its performance with each iteration.
Finally, the evaluation of the model’s performance is conducted using the test set, calculating metrics such as OA, AA, PA, and $\kappa$ to assess the quality of predictions. This comprehensive setup emphasizes the importance of hyperparameters, dataset partitioning, model architecture, and evaluation strategies in the context of HSI classification, providing a clear framework for reproducibility.
IV Experimental Datasets
This section introduces the datasets used for evaluating the proposed model, highlighting their spectral and spatial characteristics and source details. Table I overviews the class names, initial training sample sizes, validation pool, and test set samples.
The Salinas (SA) dataset is a hyperspectral remote sensing image captured by the Airborne Visible/Infrared Imaging Spectrometer (AVIRIS) sensor, depicting the Salinas Valley region in southern California, USA. This dataset consists of images measuring 512 by 217 pixels, with a spatial resolution of 3.7 meters and a spectral resolution ranging from 9.7 to 12 nm, covering a spectral range of 400 to 2500 nm. It encompasses 16 ground object categories. The original dataset includes 224 spectral bands, but 20 bands affected by atmospheric water absorption and low signal-to-noise ratio (SNR) are typically excluded, leaving 204 bands for analysis.
The Pavia University (PU) was gathered using the Reflective Optics System Imaging Spectrometer (ROSIS) sensor during an aerial campaign over Pavia in northern Italy. This dataset comprises 103 spectral bands and has a spatial resolution of 1.3 meters, with an image size of 610 $\times$ 340 pixels. Prior to analysis, spectral samples without relevant information must be discarded. The ground truth data for this HSI is divided into 9 distinct classes.
TABLE I: Training, Pool, and Test Sample Counts for Each Class Across Five Hyperspectral Datasets (WHU-Hi-HongHu (HH), WHU-Hi-HanChuan (HC), Salinas (SA), University of Houston (UH), and Pavia University (PU)). Each column group represents the class names along with the corresponding number of samples allocated to training, pooling, and testing for each dataset.
WHU-Hi-HongHu (HH) WHU-Hi-HanChuan (HC) Salinas (SA) University of Houston (UH) Pavia University (PU) Qingyun (QUH)
Class Train Pool Test Class Train Pool Test Class Train Pool Test Class Train Pool Test Class Train Pool Test Class Train Pool Test
Red roof 70 6951 7020 Strawberry 224 22143 22368 Broccoli green weeds 1 10 995 1004 Healthy grass 6 619 626 Asphalt 33 3282 3316 Trees 1391 137684 139075
Road 18 1738 1756 Cowpea 114 11262 11377 Broccoli green weeds 2 19 1844 1863 Stressed grass 6 621 627 Meadows 93 9232 9324 Concrete building 897 88859 89756
Bare soil 109 10801 10911 Soybean 51 5093 5143 Fallow 10 978 988 Synthetic grass 4 345 348 Gravel 10 1040 1049 Car 69 6822 6892
Cotton 816 80826 81643 Sorghum 27 2649 2677 Fallow rough plow 7 690 697 Trees 6 616 622 Trees 15 1517 1532 Ironhide building 49 4835 4883
Cotton firewood 31 3078 3109 Water spinach 6 594 600 Fallow smooth 13 1326 1339 Soil 6 615 621 Painted metal sheets 7 665 673 Plastic playground 1088 107779 108868
Rape 223 22055 22279 Watermelon 23 2244 2266 Stubble 20 1959 1980 Water 2 161 162 Bare Soil 25 2490 2514 Asphalt road 1280 126693 127973
Chinese cabbage 120 11931 12052 Greens 29 2922 2952 Celery 18 1771 1790 Residential 7 627 634 Bitumen 7 658 665 Total 4774 472672 477447
Pakchoi 20 2007 2027 Trees 90 8899 8989 Grapes untrained 56 5579 5636 Commercial 6 616 622 Self-Blocking Bricks 18 1823 1841
Cabbage 54 5356 5409 Grass 47 4688 4734 Soil vineyard develop 31 3070 3102 Road 6 620 626 Shadows 5 468 474
Tuber mustard 62 6135 6197 Red roof 52 5206 5258 Corn senesced green weeds 16 1623 1639 Highway 6 607 614 Total 213 21175 21388
Brassica parachinensis 55 5453 5507 Gray roof 84 8372 8455 Lettuce romaine 4wk 5 529 534 Railway 6 611 618
Brassica chinensis 45 4432 4477 Plastic 18 1821 1840 Lettuce romaine 5wk 10 954 963 Parking Lot 1 6 610 617
Small Brassica chinensis 113 11140 11254 Bare soil 46 4512 4558 Lettuce romaine 6wk 5 453 458 Parking Lot 2 3 232 234
Lactuca sativa 37 3641 3678 Road 93 9187 9280 Lettuce romaine 7wk 5 530 535 Tennis Court 2 212 214
Celtuce 5 496 501 Bright object 6 562 568 Vineyard untrained 36 3598 3634 Running Track 3 327 330
Film covered lettuce 36 3595 3631 Water 377 37324 37700 Vineyard vertical trellis 9 895 903 Total 75 7439 7515
Romaine lettuce 15 1490 1505 Total 1287 127478 128765 Total 270 26794 27065
Carrot 16 1593 1608
White radish 44 4312 4356
Garlic sprout 17 1726 1743
Broad bean 7 657 664
Tree 20 2000 2020
Total 1933 191413 193347
The WHU-Hi-HanChuan (HC) dataset was acquired on June 17, 2016, between 17:57 and 18:46 in Hanchuan, Hubei Province, China. The data collection utilized a Headwall Nano-HSI sensor with a 17-mm focal length, mounted on a Leica Aibot X6 UAV V1 platform. The acquisition occurred under clear skies, with a temperature of around $30^{\circ}\text{C}$ and relative humidity of approximately 70%. This dataset encompasses a rural-urban transitional zone that includes various features such as buildings, water bodies, and agricultural fields. Specifically, it captures seven crop types: strawberry, cowpea, soybean, sorghum, water spinach, watermelon, and leafy greens. The UAV operated at an altitude of 250 meters, producing images with dimensions of 1217 $\times$ 303 pixels across 274 spectral bands, covering a range from 400 to 1000 nm. The spatial resolution of the imagery is approximately 0.109 meters. It is important to note that the dataset contains substantial shadowed areas due to the low solar elevation angle during the late afternoon acquisition.
The WHU-Hi-HongHu (HH) dataset was collected on November 20, 2017, between 16:23 and 17:37 in Honghu City, Hubei Province, China, using a Headwall Nano-Hyperspec imaging sensor with a 17-mm focal length mounted on a DJI Matrice 600 Pro UAV. The data acquisition occurred under cloudy conditions, with an ambient temperature of approximately 8°C and a relative humidity of around 55%. This dataset covers a diverse agricultural landscape with multiple crop types, including different cultivars of the same species, such as Chinese cabbage and cabbage, as well as Brassica chinensis and small Brassica chinensis. The UAV operated at an altitude of 100 meters, capturing imagery with dimensions of 940 $\times$ 475 pixels across 270 spectral bands, spanning wavelengths from 400 to 1000 nm, and providing a spatial resolution of roughly 0.043 meters.
The University of Houston (UH) dataset was collected during the 2013 IEEE GRSS Data Fusion Contest and has been extensively utilized due to its diverse land cover types and challenging classification scenarios. It consists of hyperspectral imagery captured by the Compact Airborne Spectrographic Imager (CASI), which records reflectance data across 144 spectral bands, ranging from 380 nm to 1050 nm. The dataset provides a rich spectral profile for each pixel, enabling detailed analysis of the spatial-spectral characteristics of various land cover classes. The UH covers a spatial region of 349 $\times$ 1905 pixels with a spatial resolution of 2.5 meters per pixel, representing an area in the UH campus and its surrounding urban landscape. The region includes complex urban and natural features such as vegetation, buildings, roads, and parking lots, making it an ideal testbed for HSI classification techniques. The dataset is challenging due to the variability in spectral signatures caused by different lighting conditions, atmospheric effects, and material diversity. Ground truth labels are available for 15 land cover classes, representing diverse categories. These classes differ significantly in their spatial extent and spectral characteristics, providing a rigorous evaluation framework for classification models.
The QUH-Qingyun dataset was collected on 18 May 2021, near Qingyun Road Primary School and the surrounding residential area in Qingdao, China. Using a UAV at an altitude of 300 meters, images were captured with a spatial resolution of approximately 0.15 meters, an image size of 880 $\times$ 1360 pixels, and 270 spectral bands covering wavelengths from 400 to 1000 nm.
V Active Transfer Learning Effects on the same dataset
The results presented in Table II highlight the consistent improvement in classification accuracy as the number of training samples increases. Most classes show steady progress, with notable gains in challenging categories such as Water, which improves from 49.4% to 98.1%, and Residential, which rises from 57.9% to 94.5%. These trends demonstrate the model’s ability to effectively utilize additional training data to improve performance, particularly for underrepresented classes. Certain classes, such as Soil and Running Track, achieve near-perfect accuracy with minimal training samples, suggesting they are well-captured in the feature space. However, difficult classes like Parking Lot 2 begin with very low accuracy (1.7%) but see a significant improvement, reaching 77.8% as the sample size increases. This emphasizes the importance of ensuring sufficient representation of all classes in the training set.
@@ V Active Transfer Learning Effects on the same dataset
TABLE II: UH Dataset: Per-class accuracy for training samples, showing accuracy improvements as the sample count increases per iteration. Additionally, average training and testing times (in seconds) and the total FLOPs and parameter count provide a comprehensive overview of computational requirements.
Class Per-Class Accuracy for Training Samples
@@
Healthy grass 86.4217 95.5271 96.3258 97.2843 98.0830 98.4025
Stressed grass 97.4481 99.5215 98.8835 97.9266 99.0430 99.5215
Synthetic grass 93.1034 95.1149 91.0919 93.1034 98.2758 100
Trees 95.0160 97.2668 94.3729 97.1061 99.1961 97.9099
Soil 100 99.8389 100 100 100 100
Water 49.3827 83.9506 90.7407 93.8271 95.6790 98.1481
Residential 57.8864 73.1861 90.2208 93.8485 94.4794 94.4794
Commercial 62.3794 61.2540 83.9228 94.0514 95.1768 95.9807
Road 60.8626 90.5750 96.3258 96.1661 96.4856 97.9233
Highway 93.3224 83.3876 88.9250 98.5342 98.8599 100
Railway 35.5987 95.1456 92.8802 96.4401 98.8673 99.0291
Parking Lot 1 53.8087 75.8508 85.0891 91.0858 91.0858 97.5688
Parking Lot 2 1.7094 23.0769 46.5811 66.6666 68.8034 77.7777
Tennis Court 85.0467 91.1214 100 98.1308 99.5327 100
Running Track 93.3333 100 100 100 100 97.8787
@@
[…]
Figure 2: UH Dataset: OA, AA, and $\kappa$ accuracy as a function of training samples in each iteration as shown in Table II.
@@
Table II also provides computational metrics, showcasing the model’s efficiency. Training times remain stable at around 44 seconds across iterations, despite the increase in sample size, indicating scalability. The testing time is consistently around 3.4 seconds, demonstrating the method’s suitability for real-time applications. The reported FLOPs (110,336) and parameter count (836,559) reflect the model’s balance between computational efficiency and capacity, making it practical for resource-constrained environments. Figure 2 further supports these findings by illustrating the trends in accuracy metrics—OA, AA, and $\kappa$—as the number of training samples increases. All metrics exhibit significant growth, with the most substantial improvements occurring at lower sample counts (75 to 223), followed by a diminishing rate of improvement beyond 510 samples. OA consistently outperforms AA and $\kappa$, indicating robust performance across most classes while maintaining reliability. The steady upward trend demonstrates the model’s ability to leverage additional training data effectively.
Qualitative improvements are visually evident in Figure 3, which presents ground truth maps for varying training sample sizes. With fewer training samples (e.g., 75 and 223), the maps exhibit substantial noise and misclassification. As the training sample size increases, the maps become progressively refined, reflecting improved spatial and spectral consistency. Classes such as Road and Water demonstrate better spatial coherence at higher sample sizes (650 and 787), while sparse classes like Parking Lot 2 clearly reduce misclassification.
The results presented in Table III and Figures 4 and 5 collectively highlight the model’s performance progression on the PU dataset as training samples increase. Table III provides a detailed per-class accuracy breakdown, showing substantial accuracy improvements for most classes as sample counts rise, especially for challenging classes like Gravel and Shadows, which see increases of over 20 percentage points. This indicates that the model’s learning efficacy improves with a larger training dataset, supporting robust classification for diverse material types. Notably, Metal Sheets achieves perfect classification (100%) consistently as sample size grows. Training and testing times remain relatively stable across sample sizes, indicating computational efficiency; a total of 108,800 FLOPs and 835,017 parameters suggests a well-optimized model that balances accuracy and computational overhead. The accompanying Figure further substantiates these findings by plotting OA, AA, and $\kappa$ scores against sample counts, illustrating a clear upward trend with convergence near 99% as sample sizes approach 2245. This consistent improvement across all three metrics emphasizes that the model achieves high precision and reliability across iterations, which is critical for effective hyperspectral image classification applications in remote sensing.
TABLE III: PU Dataset: Per-class accuracy for training samples, showing accuracy improvements as the sample count increases per iteration. Additionally, average training and testing times (in seconds) and the total FLOPs and parameter count provide a comprehensive overview of computational requirements.
Class Per-Class Accuracy for Training Samples
@@
Asphalt 87.7563 96.8938 96.5319 96.8938 98.5524 98.7937
Meadows 96.9540 98.3805 99.3350 99.7211 99.7640 99.9034
Gravel 69.5900 82.1734 80.8388 84.9380 87.8932 91.3250
Trees 84.3342 91.8407 95.1697 98.1070 98.6292 99.5430
Metal Sheets 99.8514 99.5542 100 100 100 100
Bare Soil 72.7128 80.7080 97.0564 98.9260 99.3237 99.6420
Bitumen 70.0751 87.5187 97.8947 96.2406 99.3984 99.5488
Bricks 75.0678 82.1835 92.0695 94.0249 94.7311 95.8718
Shadows 66.0337 92.6160 97.0464 98.3122 97.8902 98.9451
@@
[…]
Figure 4: PU Dataset: OA, AA, and $\kappa$ accuracy as a function of training samples in each iteration as shown in Table III.
@@
Table IV and Figures 6 and 7 provide a thorough analysis of the model’s performance, computational complexity, and the scalability of its classification accuracy on the SA dataset. Table IV presents per-class accuracy for various sample sizes, progressively increasing from 270 to 2842 samples. The results highlight the model’s robust performance across most classes, achieving near-perfect accuracy as the number of training samples increases, particularly for classes like Broccoli 1, Celery, and Stubble. However, for more challenging classes, such as Grapes and Vinyard Untrained, a noticeable improvement in accuracy is observed with larger sample sizes. This demonstrates the model’s need for a substantial amount of data for better performance in complex classifications. Additionally, the table includes average training and testing times to emphasize computational efficiency. Although training times gradually increase as sample sizes grow, testing times remain relatively stable, underscoring the model’s scalability and efficiency. Furthermore, Table IV provides computational metrics, including FLOPs and parameter counts, which quantify the model’s complexity at approximately 110,592 FLOPs and 836,816 parameters. These values confirm the model’s high computational efficiency, which is an essential consideration for large-scale applications.
TABLE IV: SA Dataset: Per-class accuracy for training samples, showing accuracy improvements as the sample count increases per iteration. Additionally, average training and testing times (in seconds) and the total FLOPs and parameter count provide a comprehensive overview of computational requirements.
Class Per-Class Accuracy for Training Samples
@@
Broccoli 1 95.9163 100 100 100 99.5019 100
Broccoli 2 99.9463 100 100 100 100 100
Fallow 99.6963 98.4817 100 100 100 100
Fallow Rough 97.8479 98.5652 100 99.7130 100 100
Fallow Smooth 95.9671 98.2823 97.7595 99.4772 99.4025 99.4025
Stubble 99.4444 99.8484 100 100 100 100
Celery 100 99.4972 99.8882 99.9441 100 100
Grapes 84.6344 94.1447 97.0546 97.4627 98.2966 98.8644
Soil Vinyard 100 100 99.9677 100 100 99.9677
Corn Senesced 94.6918 98.6577 99.8169 100 100 99.9389
Lettuce 4wk 90.4494 100 100 99.4382 100 100
Lettuce 5wk 100 100 100 100 100 100
Lettuce 6wk 97.8165 99.3449 100 99.5633 100 100
Lettuce 7wk 97.7570 97.1962 99.8130 99.8130 99.8130 99.8130
Vinyard Untrained 75.8117 89.1304 92.1849 95.1568 96.9179 98.2388
Vinyard Vertical 97.5636 99.7785 100 100 100 100
@@
[…]
Figure 6: SA Dataset: OA, AA, and $\kappa$ accuracy as a function of training samples in each iteration as shown in Table IV.
@@
Figure 6 offers a visual representation of the model’s accuracy improvements with increased training samples by plotting OA, AA, and $\kappa$. All metrics display an upward trend, achieving near-perfect accuracy as the sample size approaches 2842, demonstrating that the model effectively learns and generalizes as it is exposed to more data. Notably, the close convergence of OA, AA, and $\kappa$ toward 100% signifies the model’s balanced accuracy across metrics and low classification error when trained on sufficient data. Finally, Figure 7 provides ground truth maps generated with different training sample sizes, visually showcasing classification performance improvements. The maps show less distinct class boundaries with smaller sample sizes, while with larger sample sizes, the maps display sharper and more accurate distinctions between classes. This visual progression emphasizes the model’s enhanced ability to correctly classify and delineate classes as it trains on more data.
Table V provides a comprehensive overview of the HC dataset’s classification performance and computational demands across six iterations with increasing training sample sizes. The per-class accuracy shows a consistent improvement for all classes as more samples are added, emphasizing the importance of larger datasets for enhancing model learning. Classes like Strawberry and Sorghum achieve near-perfect accuracy early, suggesting distinct and easily separable spectral features. Conversely, classes such as Plastic and Bright Object, which start with low accuracy, exhibit significant improvements, indicating the model’s ability to learn complex patterns with additional training data. The computational metrics reveal a gradual increase in training time with more samples, reflecting the added processing requirements, while testing times remain relatively stable. Furthermore, the reported FLOPs and parameter counts provide insight into the model’s computational efficiency, balancing accuracy with resource usage. Overall, the table highlights the scalability and effectiveness of the model in handling diverse classes and increasing sample sizes.
TABLE V: HC Dataset: Per-class accuracy for training samples, showing accuracy improvements as the sample count increases per iteration. Additionally, average training and testing times (in seconds) and the total FLOPs and parameter count provide a comprehensive overview of computational requirements.
Class Per-Class Accuracy for Training Samples

### v2

@@ III-A Spectral-Spectral Transformer (SST)
where $\mathbf{Q}_{i}=\mathbf{Z}_{i}\mathbf{W}_{Q}$, $\mathbf{K}_{i}=\mathbf{Z}_{i}\mathbf{W}_{K}$, $\mathbf{V}_{i}=\mathbf{Z}_{i}\mathbf{W}_{V}$, and $d_{k}$ is the attention key dimension. The output of one encoder block is:
@@
where $\mathbf{F}$ is a feedforward network:
$\mathbf{F}(\mathbf{Z}_{i})=\text{ReLU}(\mathbf{Z}_{i}\mathbf{W}_{1}+b_{1})\mathbf{W}_{2}+b_{2}$ (8)
To improve robustness against domain-specific spectral distortions, we introduce uncertainty-aware attention scaling. Let $\mathcal{U}_{i}$ be the entropy-based uncertainty for the $i$-th token. The calibrated attention is given by:
$\mathbf{A}_{\text{cal}}=\text{Softmax}\left(\frac{\mathbf{QK}^{T}}{\sqrt{d_{k}}}\right)\cdot(1+\lambda\cdot\mathcal{U})$ (9)
where $\lambda$ is a hyperparameter controlling uncertainty influence. This guides the model to focus more on spatial-spectral tokens with higher ambiguity, leading to better adaptation. To inject global context, a learnable token $\mathbf{Q}_{c}\in\mathbb{R}^{d}$ attends to the encoded patches:
@@
Classification Head
The aggregated features pass through a two-layer MLP with softmax:
@@
where $\mathbf{O}_{\text{final}}\in\mathbb{R}^{C}$ and $C$ is the number of classes.
III-B Active Learning
Hybrid Query Strategy
We define $X_{\text{pool}}$ as the set of samples and query_size as the number of samples to be queried in each AL round. Our strategy combines:
  • Uncertainty Sampling:
@@ III-B Active Learning
  • Diversity Sampling:
@@
    where
    $\text{Diversity}(S)=\frac{1}{m(m-1)}\sum_{j=1}^{m}\sum_{\begin{subarray}{c}k=1\\ k\neq j\end{subarray}}^{m}d_{jk}$ (15)
    and $d_{jk}$ is the Euclidean distance between pixel spectra in local neighborhoods of size $n_{\text{neighborhood}}^{2}$.
Query Index Selection
For each candidate $\mathbf{x}_{(h,w)}$, we compute:
@@
The top-query_size entries with the highest diversity are selected:
@@
where
$\pi=\text{argsort}(d)$ (18)
III-C Dynamic Layer Freezing
To facilitate efficient domain adaptation, we introduce a dynamic layer freezing policy. During transfer to a new target domain, SST layers are selectively frozen or unfrozen based on sensitivity to domain shift. We estimate shift magnitude using the maximum mean discrepancy (MMD) of intermediate features. Layers with low variance across domains are frozen, reducing computational cost while retaining useful knowledge.
III-D Model Training and Evaluation
The model is trained using categorical cross-entropy:
@@ III-D Model Training and Evaluation
Evaluation metrics include OA, AA, and the kappa coefficient:
@@
where $p_{o}$ is the observed agreement and $p_{e}$ is expected agreement by chance.
IV Experimental Settings and Datasets
The experimental configuration for SST-ATL is designed to ensure clarity, reproducibility, and robust performance evaluation. Following the preprocessing of the HSI cubes and corresponding ground truths, patch-based datasets are generated. Each HSI is divided into overlapping spatial windows of size $8\times 8$ (i.e., input patch size), which serve as local contextual inputs for the SST backbone. The data are partitioned into training, pool, and test sets with a split ratio of 1%, 49%, and 50%, respectively. Specifically, 1% of the samples are randomly selected for initial training, 49% are allocated to the pool for AL queries, and 50% are reserved exclusively for testing. For diversity-based query selection, a neighborhood size of $n_{\text{neighborhood}}=3$ is used, resulting in $9$ spectral vectors per neighborhood to compute diversity scores, and the query percentage is set to 0.02%.
@@ IV Experimental Settings and Datasets
Model training is conducted under a set of carefully tuned hyperparameters. The learning rate is initialized at 0.001 with a decay factor of $1\times 10^{-6}$, and the Adam optimizer is employed to facilitate convergence. Training proceeds for 50 epochs with a batch size of 56. The backbone architecture comprises four transformer layers, each featuring eight attention heads. The model dimensionality ($d_{\text{model}}$) is set to 54, and the feedforward network dimensionality is configured as $4\times 64=256$. A dropout rate of 0.1 is applied to mitigate overfitting, while layer normalization is stabilized with an epsilon value of $1\times 10^{-6}$.
The experimental evaluation in this study was conducted using six widely recognized hyperspectral datasets: Qingyun (QUH), WHU-Hi-HongHu (HH), WHU-Hi-HanChuan (HC), Salinas (SA), University of Houston (UH), and Pavia University (PU). Each dataset encompasses a variety of land cover classes with distinct characteristics, enabling a comprehensive assessment across diverse scenes. For HH and HC, a rich set of agricultural and urban classes was included, while SA predominantly featured vegetation-related categories. UH captured a complex urban environment, and PU primarily focused on urban materials and structures. The QUH dataset significantly expanded the scale and diversity of the evaluation, offering a substantial number of samples across natural and man-made categories.
V ATL-SST Effects on the same dataset
The results presented in Table II highlight the consistent improvement in classification accuracy as the number of training samples increases. Table II also provides computational metrics, showcasing the model’s efficiency. Training times remain stable at around 44s across iterations, despite the increase in sample size, indicating scalability. The testing time is consistently around 3.4s, demonstrating the method’s suitability for real-time applications. The reported FLOPs (110,336) and parameter count (836,559) reflect the model’s balance between computational efficiency and capacity, making it practical. Qualitative improvements are visually evident in Figure 2, which presents ground truth maps for varying training sample sizes. With fewer training samples (e.g., 75 and 223), the maps exhibit substantial noise and misclassification. As the training sample size increases, the maps become progressively refined, reflecting improved spatial and spectral consistency.
@@ V ATL-SST Effects on the same dataset
TABLE II: UH Dataset: Accuracy improvements as the sample count increases per iteration. Additionally, average training and testing times (in seconds) and the total FLOPs and parameter count provide a comprehensive overview of computational requirements.
Class Number of Training Samples
@@
$\kappa$ 71.27 84.98 91.00 94.96 96.17 97.38
OA 73.49 86.13 91.68 95.34 96.46 97.57
AA 71.02 84.32 90.35 94.27 95.57 96.97
@@
The results presented in Table III and 3 collectively highlight the model’s performance progression on the PU dataset as training samples increase. Training and testing times remain relatively stable across sample sizes, indicating computational efficiency; a total of 108,800 FLOPs and 835,017 parameters suggests a well-optimized model that balances accuracy and computational overhead. The accompanying Figure further substantiates these findings by plotting OA, AA, and $\kappa$ scores against sample counts, illustrating a clear upward trend with convergence near 99% as sample sizes approach 2245.
TABLE III: PU Dataset: Accuracy improvements as the sample count increases per iteration.
Class Number of Training Samples
@@
$\kappa$ 82.69 90.62 95.65 96.98 97.85 98.51
OA 87.11 92.98 96.72 97.72 98.38 98.87
AA 80.26 90.20 95.10 96.35 97.35 98.17
@@
Table IV and 4 thoroughly analyze the model’s performance, computational complexity, and the scalability of its classification accuracy on the SA dataset. Although training times gradually increase as sample sizes grow, testing times remain relatively stable, underscoring the model’s scalability and efficiency. Furthermore, Table IV provides computational metrics, including FLOPs and parameter counts, which quantify the model’s complexity at approximately 110,592 FLOPs and 836,816 parameters. These values confirm the model’s high computational efficiency, which is an essential consideration for large-scale applications. Finally, Figure 4 provides ground truth maps generated with different training sample sizes, visually showcasing classification performance improvements.
TABLE IV: SA Dataset: Accuracy improvements as the sample count increases per iteration.
Class Number of Training Samples
@@
$\kappa$ 91.55 96.59 97.99 98.62 99.08 99.42
OA 92.41 96.94 98.20 98.76 99.17 99.48
AA 95.47 98.30 99.15 99.41 99.62 99.76
@@
Table V provides a comprehensive overview of the HC dataset’s classification performance and computational cost. The computational metrics reveal a gradual increase in training time with more samples, reflecting the added processing requirements, while testing times remain relatively stable. Furthermore, the reported FLOPs and parameter counts provide insight into the model’s computational efficiency, balancing accuracy with resource usage. Figure 5 showcases the ground truth maps for each training sample size, visually representing the model’s learning progress. With 1,287 samples, the classification appears fragmented, reflecting limited learning. As the sample size increases, the maps exhibit progressively finer details, and class boundaries become more accurate and well-defined.
TABLE V: HC Dataset: Accuracy improvements as the sample count increases per iteration.
Class Number of Training Samples